OpenStudy (anonymous):

A 59 kg man stands still in an elevator moving with a constant speed of 1.9 m/s. a) if the down going elevator comes to rest with a constant deceleration over a distance of 4.2 m, what is the magnitude of the force the elevator floor exerts on the man? b) if the elevator is moving up before the deceleration, with the other conditions the same as in part (a), what is the magnitude of that force now?

OpenStudy (anonymous):

\[V^2 - u^2 = 2as\] \[\frac{ v^2 - u^2 }{ 2s } = a\]